2. Find the dy/dt for each of the following: a) y = 24 b) y = te c) y=teat d) y = log(7) e) y = log|7t(1 - )
Answer #1

2 a) dy/dt = 2e2t+4

b) dy/dt = tet + et

c) dy/dt = 2t2e2t + 2te2t

d) dy/dt = (1/ 7t5)* 35t4 = 5/t

e) dy/dt = (7(1−t)5−35(1−t)4*t)/ (7(1−t)5*t) = (6t-1) / ((t-1)*t)